Subject: 3.5-m Users Committee minutes 1/22/07

From: Bruce Gillespie

Submitted: Tue, 23 Jan 2007 18:40:12 -0700

Message number: 1009 (previous: 1008, next: 1010 up: Index)

                         APO 3.5-m Users Committee Phonecon, 1/22/07

Attending:  Suzanne Hawley, Bruce Gillespie, Michael Strauss, Scott  
Anderson, Jon Holtzman, Russet McMillan, John Bally

Absent: Don York, Rene Walterbos, Remy Indebetouw, John Fulbright, Al  

Minutes taken by Bruce Gillespie


User feedback, comments from institutional representatives:

Princeton (Strauss) - Michael said that the PU users are wishing for  
better weather.  There is some frustration users have had with the  
user documentation at the APO website.  Suzanne mentioned that  
Gabrelle Saurage has been enlisted to produce a basic users manual  
for each instrument, and is working these with help from Jon Holtzman  
and Jack Dembicky.  The DIS manual is ready to go, and users are  
invited to go look at it and give us comments; see http:// 

JHU (Fulbright) - no report

Chicago (Harper) - no report

Washington (Anderson) - Scott said that there had been UW users  
concerned with the DIS cal lamps issue (see telescope report) and  
with the problems with the DIS slit view camera shutter.

NMSU (Holtzman) - Jon said that "stray light" has recently been seen  
in DIS images.  We're trying to confirm that it is caused by the CCD  
heater, which was recently turned back on when we installed the new  
red CCD.

Virginia (Indebetouw) - no report

Colorado (Bally) - John reported that he's recently used DIS with the  
new red CCD, and that he's seen that the fringing is greatly improved.


Discussion of telescope/instruments report:

The report is given below.


      3.5-m Telescope, Instruments, and CIF Projects Highlights,  
11/23/06 through 1/18/07

                                  Bruce Gillespie

0) Overview

Winter weather has returned to the area, and the associated  
cloudiness has frustrated several observing programs.  But whenever  
conditions have allowed, the staff and equipment were able to eke out  
good science data.  The good highlight for the period was the  
successful installation of the new deep-depletion CCD for the red  
camera of DIS--the new chip appears to have good sensitivity and much  
reduced fringing in the deep red spectrum.  The biggest trouble we  
have had is an unreliable shutter in the DIS slitviewer.

1) Telescope

Other than routine maintenance, not much happened with the telescope  
during the holidays, and it performed very well.  Some tuning of the  
drive servos was done to improve tracking performance, and a new  
version of TUI is being site-tested prior to release to the users.

2) Instruments

DIS:  Now that the DIS upgrade project is substantially complete, we  
have found that the improved sensitivity and performance at both ends  
of the spectrum has uncovered shortcomings in our lamp calibration  
system.  Our arc lamps have inadequate lines (mainly in the UV) and  
the flat-field lamps have inadequate intensity.  We also will need  
decent arc lamps for TripleSpec later this year.  We have decided to  
undertake an overall study of the lamp calibration system  
requirements for the telescope and all instruments, and plan to  
invest in some new lamps, optics, and flat-field reflecting surface  
(on the mirror covers) during the course of the year.

Throughput measurements of DIS were made, showing substantial  
improvement in both the UV and red parts of the spectrum.  The new  
R1200 grating was damaged during handling while the grating tilt was  
being adjusted.  We have kludge-reinstalled the grating, but are  
going to replace it anyway.  The DIS slitviewer camera shutter became  
balky during the cold weather.  We have replaced the shutter and  
modified the ventilation around the camera to keep it warmer during  
the winter, but are also looking into shutterless operation of the  

NIC-FPS:  Stephane Beland is coming to APO next week to look into  
recently reported problems with the camera controller.

3) Engineering and CIF projects

Work continues on TripleSpec, heading for a delivery to the site  
possibly this summer.  A plan to modify the NA2 instrument rotator is  
underway, and a design requirements discussion was held for a future  
CIF project to rebuild the tertiary mount and actuators.  We have  
also started work on upgrading the telescope drive controllers to  
make them perform better and be maintainable.

4) Miscellaneous

Nick McDonald began work as the new 3.5-m Telescope Systems Engineer,  
refilling the post Jon Davis vacated last year.


Discussion of new site software protocols - group:

Suzanne asked how the new password and security protocols were  
working out for the users.  Everybody said they were OK.  Scott said  
that he was aware of problems people were having ftp-ing their data  
from APO, some kind of timeout problem breaking the connection.  It  
wasn't immediately clear which end of the connection was causing this  
problem, so Bruce will get Fritz to talk to Scott to try to work this  


Discussion on fast guiding, rotator upgrade - group:

Suzanne gave a short report on the NA2 instrument rotator upgrade  
project, planned for this year, which is being done in order to fix  
long-standing problems with the rotator and to give it better  
capacity to handle TripleSpec.  We plan for a PDR in early April, and  
the design work is being done by Mike Carr at Princeton.  On fast  
guiding, we are looking at this (again) partly because we're starting  
the redesign of the tertiary mirror mount and actuator system, and  
are wondering if tip-tilt provisions should be built in.  Instead of  
using the tertiary, it could be possible that a low-bandwidth tip- 
tilt capability could be implemented at the 2ndary mirror, now that  
the new top end is giving us better mechanical mount stiffness.  But  
would it be worth it?  Apocryphal reports from people at other  
telescope suggest that tip-tilt offers only modest image quality  
improvements.  We will explore this some more, and welcome any  
comments from our users on the tip-tilt issue.


Agile update - Hawley:

Agile, the high-speed photometer being built by Anjum Makadum, has  
its commissioning run scheduled for mid-March.  All the parts are in  
hand, and Russell Owen and Ed Mannery are helping with the s/w and h/ 
w.  It's possible that Agile could be available for shared-risk use  
in Q2 2007, depending on how the commissioning run goes.


Actions from earlier meetings - group

[open from previous months]:


[new actions from last month]:

ACTION:  Rene to collect history of 3.5-m publications, and Bruce  
will make a space for them on the APO website.
STATUS:  Rene and Bruce are working on this, and it may be possible  
to use the list Don York has produced for the ARC Futures document.   
This list is presently being vetted for accuracy.


Next meeting:

The next Users Committee phonecon will be on Monday, 26 February, at  
8:30 AM Pacific Time.  Agenda and other materials will be sent to the  
committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1009 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to